You want to build a rectangular garden area with a perimeter of 140 meters. You also want the width to be 40 meters more than the length. Find the length and width of your possible garden.
Answer by mananth(16949)

length be x
width will be x+40
2(x+40) + 2x= 140
2x+80 +2x= 140
4x=140-80
4x=60
x= 15
15m & 55m are the dimensions of the garden